OpenStudy (anonymous):

$7.80 to $4.52 simplify as fraction..please help..

Directrix (directrix):

7.80/4.52 = 7.80/4.52 * 100/100 = 780/452 = We have to simplify this fraction.

OpenStudy (anonymous):

ok how do i do that...

Directrix (directrix):

We have to look for common factors of 780 and 452. Both are even so 2 will divide both. Some people look for the greatest common factor right off the bat. I just start with 2 and work on down.

OpenStudy (anonymous):

39/226?

Directrix (directrix):

780 = 2*390 452 = 2* 226

Directrix (directrix):

780 = 2*390 = 2* 195 452 = 2* 226 = 2* 113

OpenStudy (anonymous):

195/113

Directrix (directrix):

195/113 I'll look up a few of the divisibility rules for you.

OpenStudy (anonymous):

ty

Directrix (directrix):

The 2, 3, 5, and 10 rules are the ones I try to remember. A person can live without them.:)